News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 23 September 2022 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

A 10 ks ACIS observation of M15X-2 on Sep 19 started the science run but did not complete the bias computation, and no event data were produced for this observation. All observations following this executed nominally, as was expected given the power cycling that ACIS executes between observations that recompute bias. The ACIS team is investigating this event and the observation of M15X-2 is being rescheduled.

Real-time procedures were executed on Sep 16 and Sep 20 to conduct a checkout of the HRC-S, using the side-A electronics and including a 15 ks observation of a celestial target. The HRC team is evaluating the data.

An additional real-time procedure was executed on Sep 21 to patch Thermal Secondary Database heater set points in preparation for a short-duration Telescope Forward Thermal Enclosure (TFTE) heater set point adjustment to verify TFTE, High Resolution Mirror Assembly, and Optical Bench Assembly temperature changes with new lower set points over a 4 hour period. The test was completed on Sep 22 and the spacecraft was returned to the current heater configuration by switching back to the Primary Thermal Database.

Further real-time procedures were executed on Sep 22 to dump OBC-A memory as a follow-up to the flight software patch on Sep 21 that updated the Thermal Secondary Database. The dump will be used to update the baseline memory images maintained on the ground.

The schedule of targets for the next week is shown below and includes an observation of M15X-2, which was accepted as a Target of Opportunity on Sep 13.

        CMZMolecularCloud      ACIS-I       Sep 26
        RAqr                   ACIS-S
        KeplerSNR              ACIS-S/HETG
        M15X-2                 ACIS-S/HETG  Sep 27
        Abell2744              ACIS-I
        SNR1987A               ACIS-S/HETG
        Radiation Belts                     Sep 28
        SDSSJ001900.30+1507    ACIS-S
        PTF11qcj               ACIS-S
        3CR14                  ACIS-S
        SDSSJ114804.00+45342   ACIS-S       Sep 29
        3CR14                  ACIS-S
        SDSSJ104846.63+44071   ACIS-S
        CMZMolecularCloud      ACIS-I
        Abell2744              ACIS-I       Sep 30
        Radiation Belts                     Oct  1
        CMZMolecularCloud      ACIS-I
        ColdECSBlankSky        HRC-S
        Abell141               ACIS-S
        V745_Sco_s1            ACIS-S       Oct  2
        GDH29                  ACIS-I
        SDSSJ234019.82-00590   ACIS-S
        GDH29 (4 obs)          ACIS-I
        SDSSJ234012.69+00243   ACIS-S
        CMZMolecularCloud      ACIS-I
        Radiation Belts                     Oct  3


All spacecraft subsystems except HRC continued to support nominal operations.

Return to Status Reports